# Retrieve, Merge, Predict: Augmenting Tables with Data Lakes

### v6

@@ Abstract
Machine-learning from a disparate set of tables, a data lake, requires assembling features by merging and aggregating tables. Data discovery can extend autoML to data tables by automating these steps. We present an in-depth analysis of such automated table augmentation for machine learning tasks, analyzing different methods for the three main steps: retrieving joinable tables, merging information, and predicting with the resultant table. We use two data lakes: Open Data US, a well-referenced real data lake, and a novel semi-synthetic dataset, YADL (Yet Another Data Lake), which we developed as a tool for benchmarking this data discovery task. Systematic exploration on both lakes outlines 1) the importance of accurately retrieving join candidates, 2) the efficiency of simple merging methods, and 3) the resilience of tree-based learners to noisy conditions. Our experimental environment is easily reproducible and based on open data, to foster more research on feature engineering, autoML, and learning in data lakes.
@@ 1 Introduction: learning on data lakes needs feature engineering
New tools keep facilitating data science with machine learning, partly automating it, as with autoML (He et al., 2021; Karmaker et al., 2021; Erickson et al., 2020; Feurer et al., 2022; Hutter et al., 2019). However, they typically take a single table as an input, while data scientists often start from a data lake: a loose corpora of tables (Nargesian et al., 2019; Spotify, 2020; Phan, 2023). Learning then requires assembling multiple tables, which needs either in-depth knowledge of the data or data discovery (Fan et al., 2023a; Hulsebos et al., 2024). Data assembly is typically studied in scenarios where the schema is known (Robinson et al., 2024; Wang et al., 2024): how columns of various tables are related and which can be merged, as in a relational database. However, data lakes often come without this information. In this work, we consider a data-lake situation where the meaningful merges are unknown and where only a fraction of the tables in the data lake are relevant for the given machine learning (ML) task. Consider the following example scenario:
@@
Alice, a data scientist, wants to predict the rating given by users to a movie: she has access to a table with information about movies (e.g., year of release, director, language, budget, …). She wants to complete the table, finding more information on the subject beneficial to her task; for example, joining the table about movies with a table on the cast of those movies might help, as some actors tend to appear in movies with better ratings. Alice has access to some large repository of data, or to some search engine that she can query to find additional tables (Castelo et al., 2021; Google, 2023). However, she does not know the schema describing the relation between tables in the data repository. Her objective is to find the tables that are most relevant to her task, to merge them with the original table and finally to use the improved table to build a model that predicts the movie rating. She has limited resource (compute, RAM…), and thus cannot evaluate all potential tables.
@@
Figure 1: The evaluation pipeline. Given a base table, the three main steps (Retrieve, Merge, Predict) augment it with the information from the lake to improve the prediction performance.
@@
Research on this literature is spread across the database and machine learning communities. The corresponding pipeline (Figure 1) uses retrieve, merge, and predict steps that involve four tasks: 1) retrieving tables that are joinable with the original table (Fernandez et al., 2019; Zhu et al., 2019; 2016; Castelo et al., 2021; Fernandez et al., 2018; Dong et al., 2023), 2) selecting which joins should be executed to improve the performance of the subsequent ML model (Esmailoghli et al., 2021; Liu et al., 2022; Galhotra et al., 2023; Deng et al., 2017a), 3) aggregating results in cases of one-to-many or many-to-many joins (Chepurko et al., 2020; Kanter & Veeramachaneni, 2015), 4) predicting with supervised learning on the resulting table, a tabular learning problem (Shwartz-Ziv & Armon, 2022; Borisov et al., 2022; Grinsztajn et al., 2022). Elaborate methods have been designed for each of these tasks separately. Yet, how they contribute to the overall machine learning pipeline is often unclear. Prediction performance is not evaluated consistently in the database literature. Moreover, evaluation on data lakes is particularly challenging. First, there is no openly-available reference data suited for this purpose: both a data lake and tables with corresponding analytic tasks. Second, a solid evaluation must run the pipeline many times across loosely structured, messy ensembles of tables. State-of-the-art publications relevant to this pipeline seldom come with implementations robust and scalable enough for cross-validation loops on a data lake. Indeed, data discovery and assembly code can involve many complex elements, such as probabilistic data structures or large language models, which need careful software engineering to be production ready. The difficulty of operating those complex pipelines does beg the question of when the prediction gain is worth the operational cost. State-of-the-art research often leaves aside the question of complexity, which involves not only computational cost, but also preparation and maintenance burden (Sculley et al., 2015; Paleyes et al., 2022).
@@
We study the full data-discovery for machine learning pipelines, with an eye on exploring the complexity gradient in a reproducible way. For this, we contribute Yet Another Data Lake (YADL), a dataset that enables systematical exploration of the aspects of the pipeline important for good prediction. Built from the YAGO knowledge base (Mahdisoltani et al., 2014), YADL provides a controlled environment to test methods. To analyze the discovery and merge pipelines across different degrees of complexity, YADL’s synthesis procedure allows to vary the shape (number of rows and columns) and degree of redundancy of the resulting tables.
@@
We investigate the impact of the main steps on the end goal of learning from data lakes: generating potential augmentation candidates, integrating them with the base table, and subsequently evaluating learning on this assembled data. Considering both retrieving and combining tables in a data lake and the downstream learning task, our work bridges database search and machine learning, a topic of much interest (Fey et al., ; Robinson et al., 2024; Cvetkov-Iliev et al., 2023; Wang et al., 2024). It has two main objectives: first, determine which pipeline steps are most important for prediction performance, and second, develop an equitable playing field that facilitates a holistic assessment of learning on data lakes.
@@
We analyze various methods in the four tasks, with an exhaustive empirical evaluation that required approximately 21 years or 189k CPU and GPU hours (Table 7). Our findings show that 1) automated selection of joins is noisy, creating very messy tables for which tree-based methods are more robust than deep learning models; 2) better retrieval of joins reduces the amount of noise and the fraction of null values; 3) relying on Jaccard containment (i.e., the fraction of entities of the table to augment found in a join candidate) as a criterion for retrieving joins improves downstream prediction performance; 4) finally, simpler, metric-based join-discovery methods outpace their complex counterparts in terms of speed and the larger cost of more elaborate methods does not yield proportionate gains in prediction performance.
@@ 2 Problem setting and related work
Problem setting. Consider a user training a ML model to predict some quantity. The corresponding quantity appears in the training data as a target column $Y$ of a base table $T$. Assume our user has also access to a large collection of tables (a data lake) $D=\{T_{1},T_{2},...,T_{m}\}$, some of which may contain additional information that can enrich the base table $T$. The data lake may be publicly available (Castelo et al., 2021; Mattmann et al., 2018; Srinivas et al., 2023; Hulsebos et al., 2023; Nargesian et al., 2018) or private (e.g., corporate data). Departing from many studies on learning on relational databases, where schemas are provided (Motl & Schulte, 2015; Fey et al., 2023; Robinson et al., 2024), we consider unstructured data lakes, i.e., data lakes without a schema to specify any PK-FK (Private Key - Foreign Key) relationship between the tables.
@@
Given a base table $T$ and a data lake $D$, the Candidate Retrieval task consists in discovering join candidates (Retrieve in Figure 1). This task looks for tables that can be considered as “candidate joins” for the given base table. Different methods have been developed for different scenario: some “dataset search engines” tackle internet-scale crawls of tables (Castelo et al., 2021; Mattmann et al., 2018); other methods are instantiated on data stored locally (Zhu et al., 2019; Fernandez et al., 2019; Zhu et al., 2016; Fernandez et al., 2018; Khatiwada et al., 2022; 2023b).
While integrating data is not a central focus of the ML literature, augmenting features via joins is recognized as key to ML (Paleyes et al., 2022; Kumar et al., 2016). Still, the blind addition of features may lead to diminishing or negative returns, which led to the development of systems that seek to integrate only correlated features (Esmailoghli et al., 2021; Santos et al., 2021). Alternative systems seek to augment $T$ by synthesising new features (Kanter & Veeramachaneni, 2015; Cvetkov-Iliev et al., 2023; Zhao & Castro Fernandez, 2022). Recent contributions focus on building graph representations of relational databases to leverage GNNs for learning across tables (Fey et al., ; Robinson et al., 2024; Wang et al., 2024). An orthogonal research direction is concerned with the problem of search and augmentation while maintaining privacy (Huang et al., 2023a; b).
Though implementations may vary, retrieval methods used to find join candidates share some similarities: 1) they involve an offline phase to either index the data lake (Zhu et al., 2016; Fernandez et al., 2019), or train a model (Fan et al., 2023b; Dong et al., 2023), and 2) they rely on similarity metrics across columns to select candidate joins. Often, this metric is Jaccard Containment (JC), defined as
@@
where $Q\in T$ is a query column in base table $T$, $C_{k}^{i}\in D$ is a candidate column in table $T_{k}$, $|Q\cap C_{k}^{i}|$ is the cardinality of the intersection between the two sets, and $|Q|$ is the cardinality of the query set itself. Intuitively, if this ratio is high, then a large fraction of $Q$ is found in $C_{k}^{i}$, suggesting that the two columns should be joined. Approximate methods can scale JC-based retrieval to very large databases (Zhu et al., 2016; Fernandez et al., 2019; 2018). Beyond JC, other metrics have been used, such as top-$k$ set similarity (Zhu et al., 2019), combinations of JC and embeddings-based similarities (Fan et al., 2023b), or other embeddings-based metrics (Cong et al., 2022; Habibi et al., 2020), such as approximate nearest neighbor search (Dong et al., 2023) on model embeddings. Methods based on Jaccard similarity are vulnerable to typos, which led to the development of methods that perform fuzzy or semantic matching such as Fan et al. (2023b); Mundra et al. (2023); Deng et al. (2017b); in this work, we focus on scenarios that do not involve fuzzy matching.
@@
To make retrieval readily useful, the Join Selection task identifies a subset of candidate joins that maximizes the prediction performance over a downstream task (Select in Figure 1). These methods use various strategies to add value to retrieval: profiling candidate joins according to various metrics (Galhotra et al., 2023; Flores et al., 2021), rules to remove joins that are not useful (Kumar et al., 2016; Shah et al., 2017), joining over sketches or coresets of the data (Wang et al., 2022; Santos et al., 2021), or executing each candidate join to find those that bring benefit (Esmailoghli et al., 2021; Chepurko et al., 2020; Liu et al., 2022; Galhotra et al., 2023; Gong et al., 2023; Dong & Oyamada, 2022). In this work, we focus on the latter, very popular, strategy. We consider ML models that operate on rows as samples, columns forming features. Augmentation enriches features while keeping the original set of samples, thus requiring a left join. A resulting challenging is that un-joined rows in the left table will lead to missing (null) values in the new columns.
@@
After the set of join candidates has been retrieved, they must be merged with the base table to augment it with their columns. Depending on the join selection strategy, aggregation may be executed during the selection process, or after the set of candidates has been chosen. In any case, it is paired with the selection procedure by the need of training the ML model on the augmented tables. To represent this coupling, we combine the two tasks in a single Merge step in Figure 1.
It is not always possible to limit joins to one-to-one relationships; often, we join on one-to-many or many-to-many relations. For example, to join a table about movies with one that contains movie ratings on column “movie title”, every movie with more than one rating is in a one-to-many relation, thus the content of all the rows in such relations gets duplicated. Sample duplication is problematic when the downstream task involves a ML or statistical analysis method, as each row corresponds to one sample. The Aggregation task bridges the result of the join with the downstream methods, combining the information contained from a potentially large number of rows into one (Aggregate in Figure 1). More precisely, given a tuple (row) $t$ in the base table $T$ that needs to be joined with $n$ tuples from a table ($A$, $B$) over $A$, the goal is to augment $t$ with attribute $B$ by selecting one value that represents the information from the $n$ joining tuples. How to select a representative value for the new attribute reminds data integration problems such as truth discovery (Dong et al., 2009) or data fusion (Bleiholder & Naumann, 2009). In this spirit, Deep Feature Synthesis (DFS) (Kanter & Veeramachaneni, 2015) takes a set of tables and a join plan to recursively aggregate replicated instances using functions such as average, median, and mode.

### v7

@@ Abstract
Machine-learning from a disparate set of tables, a data lake, requires assembling features by merging and aggregating tables. Data discovery can extend autoML to data tables by automating these steps. We present an in-depth analysis of such automated table augmentation for machine learning tasks, analyzing different methods for the three main steps: retrieving joinable tables, merging information, and predicting with the resultant table. We use two data lakes: Open Data US, a well-referenced real data lake, and a novel semi-synthetic dataset, YADL (Yet Another Data Lake), which we developed as a tool for benchmarking this data discovery task. Systematic exploration on both lakes outlines 1) the importance of accurately retrieving candidate tables to join, 2) the efficiency of simple merging methods, and 3) the resilience of tree-based learners to noisy conditions. Our experimental environment is easily reproducible and based on open data, to foster more research on feature engineering, autoML, and learning in data lakes.
@@ 1 Introduction: learning on data lakes needs feature engineering
New tools keep facilitating data science with machine learning, partly automating it, as with autoML (He et al., 2021; Karmaker et al., 2021; Erickson et al., 2020; Feurer et al., 2022; Hutter et al., 2019). However, they typically take a single table as an input, while data scientists often start from a data lake: a loose corpus of tables (Nargesian et al., 2019; Spotify, 2020; Phan, 2023). Learning then requires assembling multiple tables, which needs either in-depth knowledge of the data, or data discovery (Fan et al., 2023a; Hulsebos et al., 2024). Data assembly is typically studied in scenarios where the schema is known (Robinson et al., 2024; Wang et al., 2024): how columns of various tables are related and which can be merged, as in a relational database. However, data lakes often come without this information. In this work, we consider a data lake situation where the meaningful merges are unknown and where only a fraction of the tables in the data lake are relevant for the given machine learning (ML) task. Consider the following example scenario:
@@
Alice, a data scientist, wants to predict the rating given to a movie: she has access to a table with information about movies (e.g., year of release, director, language, budget, …). She wants to complete the table, finding more information on the subject beneficial to her task; for example, joining the table about movies with a table on the cast of those movies might help, as some actors tend to appear in movies with better ratings. Alice has access to some large repository of data, or to some search engine that she can query to find additional tables (Castelo et al., 2021; Google, 2023). However, she does not know the schema describing the relation between tables in the data repository. Her objective is to find the tables that are most relevant to her task, to merge them with the original table and finally to use the improved table to build a model that predicts the movie rating. To ensure reliable prediction results, a cross-validation setup repeats these steps over different train-test splits; as resources (compute, RAM…) are limited and cross-validation has a cost, Alice is interested in finding the optimal set of candidates to achieve the best results within a certain budget.
@@
Figure 1: The evaluation pipeline. Given a base table, the three main steps (Retrieve, Merge, Predict) augment it with the information from the lake to improve the prediction performance. The data preparation step can be executed offline, and the resulting index may be reused across different data usage instances.
@@
Research in this area spans the database and machine learning communities. The corresponding pipeline (Figure 1) uses retrieve, merge, and predict steps that involve four tasks: 1) retrieving tables that are joinable with the original table (Fernandez et al., 2019; Zhu et al., 2019; 2016; Castelo et al., 2021; Fernandez et al., 2018; Dong et al., 2023), 2) selecting which joins should be executed to improve the performance of the subsequent ML model (Esmailoghli et al., 2021; Liu et al., 2022; Galhotra et al., 2023; Deng et al., 2017a), 3) aggregating results in cases of one-to-many or many-to-many joins (Chepurko et al., 2020; Kanter & Veeramachaneni, 2015), 4) predicting with supervised learning on the resulting table, a tabular learning problem (Shwartz-Ziv & Armon, 2022; Borisov et al., 2022; Grinsztajn et al., 2022). Elaborate methods have been designed for each of these tasks separately. Yet, how they contribute to the overall machine learning pipeline is often unclear. Prediction performance is not evaluated consistently in the database literature. Moreover, evaluation on data lakes is particularly challenging. First, there is no openly available reference data suited for this purpose — both a data lake and tables with corresponding analytic tasks are needed. Second, a solid evaluation must run the pipeline many times across loosely structured, messy ensembles of tables. State-of-the-art publications relevant to this pipeline seldom come with implementations robust and scalable enough for cross-validation loops on a data lake. Indeed, data discovery and assembly code can involve many complex elements, such as probabilistic data structures or large language models, which need careful software engineering to be production ready. The difficulty of operating those complex pipelines does beg the question of when the prediction gain is worth the operational cost. State-of-the-art research often leaves aside the question of complexity, which involves not only computational cost, but also preparation and maintenance burden (Sculley et al., 2015; Paleyes et al., 2022).
@@
We study the full data-discovery for machine learning pipelines, with an eye on exploring the complexity gradient in a reproducible way. For this, we contribute Yet Another Data Lake (YADL), a dataset that enables systematic exploration of the aspects of the pipeline important for good prediction. Built from the YAGO knowledge base (Mahdisoltani et al., 2014), YADL provides a controlled environment to test methods. To analyze the discovery and merge pipelines across different degrees of complexity, YADL’s synthesis procedure allows us to vary the shape (number of rows and columns) and degree of redundancy of the resulting tables.
@@
We investigate the impact of the main steps on the end goal of learning from data lakes: preparing the data structures needed to search for candidates, retrieving and selecting potential augmentation candidates, integrating them with the base table, and subsequently evaluating learning on this assembled data. Considering both retrieving and combining tables in a data lake and the downstream learning task, our work bridges database search and machine learning, a topic of much interest (Fey et al., ; Robinson et al., 2024; Cvetkov-Iliev et al., 2023; Wang et al., 2024). It has two main objectives: first, determine which pipeline steps are most important for prediction performance, and second, develop a level playing field that facilitates a holistic assessment of learning on data lakes.
@@
We analyze various methods in the four tasks, with an exhaustive empirical evaluation that required approximately 21 years or 189k CPU and GPU hours (Table 7). Our findings show that 1) automated selection of joins is noisy, creating very messy tables for which tree-based methods are more robust than deep learning models; 2) better retrieval of joins reduces the amount of noise and the fraction of null values; 3) relying on Jaccard containment (i.e., the fraction of entities of the table to augment found in a join candidate) as a criterion for retrieving joins improves downstream prediction performance; 4) downstream performance plateaus rapidly as more tables are merged, while the resource cost increases; 5) finally, simpler, metric-based join-discovery methods outpace their complex counterparts in terms of speed and the larger cost of more elaborate methods does not yield proportionate gains in prediction performance.
@@ 2 Problem setting and related work
Problem setting. Consider a user training a ML model to predict some quantity. The corresponding quantity appears in the training data as a target column $Y$ of a base table $T$. Assume our user has also access to a large collection of tables (a data lake) $D=\{T_{1},T_{2},...,T_{m}\}$, some of which may contain additional information that can enrich the base table $T$. The data lake may be publicly available (Castelo et al., 2021; Mattmann et al., 2018; Srinivas et al., 2023; Hulsebos et al., 2023; Nargesian et al., 2018) or private (e.g., corporate data). Departing from many studies on learning on relational databases, where schemas are provided (Motl & Schulte, 2015; Fey et al., 2023; Robinson et al., 2024), we consider unstructured data lakes, i.e., data lakes without a schema to specify any PK-FK (Primary Key - Foreign Key) relationship between the tables.
@@
Furthermore, we consider a use case where this operation pipeline is wrapped in a cross-validation setup to ensure the reliability of the results: we describe this use case as “Data usage” in Figure 1.
@@
Given a base table $T$ and a data lake $D$, the Candidate Retrieval task consists in discovering join candidates (Retrieve in Figure 1). This task looks for tables that can be considered as “candidate joins” for the given base table. Different methods have been developed for different scenarios: some “dataset search engines” tackle internet-scale crawls of tables (Castelo et al., 2021; Mattmann et al., 2018); other methods are instantiated on data stored locally (Zhu et al., 2019; Fernandez et al., 2019; Zhu et al., 2016; Fernandez et al., 2018; Khatiwada et al., 2022; 2023b).
While integrating data is not a central focus of the ML literature, augmenting features via joins is recognized as key to ML (Paleyes et al., 2022; Kumar et al., 2016). Still, the blind addition of features may lead to diminishing or negative returns, which led to the development of systems that seek to integrate only correlated features (Esmailoghli et al., 2021; Santos et al., 2021). Alternative systems seek to augment $T$ by synthesizing new features (Kanter & Veeramachaneni, 2015; Cvetkov-Iliev et al., 2023; Zhao & Castro Fernandez, 2022). Recent contributions focus on building graph representations of relational databases to leverage GNNs for learning across tables (Fey et al., ; Robinson et al., 2024; Wang et al., 2024). An orthogonal research direction is concerned with the problem of search and augmentation while maintaining privacy (Huang et al., 2023a; b).
Though implementations may vary, retrieval methods used to find join candidates share some similarities: 1) they involve an offline phase to either index the data lake (Zhu et al., 2016; Fernandez et al., 2019), or train a model (Fan et al., 2023b; Dong et al., 2023), and 2) they rely on similarity metrics across columns to select candidate joins. For our experiments, we assume the user has run the offline step prior to the retrieval operation (the “data preparation” section to in our pipeline, Figure 1).
Often, the similarity metric is Jaccard Containment (JC), defined as
@@
where $Q\in T$ is a query column in base table $T$, $C_{k}^{i}\in D$ is a candidate column in table $T_{k}$, $|Q\cap C_{k}^{i}|$ is the cardinality of the intersection between the two sets, and $|Q|$ is the cardinality of the query set itself. Intuitively, if this ratio is high, then a large fraction of $Q$ is found in $C_{k}^{i}$, suggesting that the two columns should be joined. Approximate methods can scale JC-based retrieval to very large databases (Zhu et al., 2016; Fernandez et al., 2019; 2018). Beyond JC, other metrics have been used, such as top-$k$ set similarity (Zhu et al., 2019), combinations of JC and embeddings-based similarities (Fan et al., 2023b), or other embeddings-based metrics (Cong et al., 2022; Habibi et al., 2020), such as approximate nearest neighbor search (Dong et al., 2023) on model embeddings. Methods based on Jaccard similarity are vulnerable to typos, which led to the development of methods that perform fuzzy or semantic matching such as Fan et al. (2023b); Mundra et al. (2023); Deng et al. (2017b); Abedjan et al. (2015); in this work, we focus on scenarios that do not involve fuzzy matching and thus need to rely on an exact metric to ensure overlap between columns, such as JC.
@@
To make retrieval readily useful, the Join Selection task identifies a subset of candidate joins that maximizes the prediction performance over a downstream task (Select in Figure 1). These methods use various strategies to add value to retrieval: profiling candidate joins according to various metrics (Galhotra et al., 2023; Flores et al., 2021), rules to remove joins that are not useful (Kumar et al., 2016; Shah et al., 2017), joining over sketches or coresets of the data (Wang et al., 2022; Santos et al., 2021), or executing each candidate join to find those that bring benefit (Esmailoghli et al., 2021; Chepurko et al., 2020; Liu et al., 2022; Galhotra et al., 2023; Gong et al., 2023; Dong & Oyamada, 2022). In this work, we focus on the latter, very popular, strategy. We consider ML models that operate on rows as samples, columns forming features. Augmentation enriches features while keeping the original set of samples, thus requiring a left join. A challenging consequence is that un-joined rows in the left table will lead to missing (null) values in the new columns.
@@
After the set of join candidates has been retrieved, they must be merged with the base table to augment it with their columns. It is not always possible to limit joins to one-to-one relationships; often, we join on one-to-many or many-to-many relations. For example, to join a table about movies with one that contains movie ratings on column “movie title”, every movie with more than one rating is in a one-to-many relation, thus the content of all the rows in such relations gets duplicated [^1] More detail is provided in Appendix subsection B.3. Sample duplication is problematic when the downstream task involves a ML or statistical analysis method, as each row corresponds to one sample.
The Aggregation task bridges the result of the join with the downstream methods, combining the information contained from a potentially large number of rows into one (Aggregate in Figure 1). More precisely, given a tuple (row) $t$ in the base table $T$ that needs to be joined with $n$ tuples from a table ($A$, $B$) over $A$, the goal is to augment $t$ with attribute $B$ by selecting one value that represents the information from the $n$ joining tuples. How to select a representative value for the new attribute reminds data integration problems such as truth discovery (Dong et al., 2009) or data fusion (Bleiholder & Naumann, 2009). In this spirit, Deep Feature Synthesis (DFS) (Kanter & Veeramachaneni, 2015) takes a set of tables and a join plan to recursively aggregate replicated instances using functions such as average, median, and mode.
Depending on the join selection strategy, aggregation may be executed during the selection process, or after the set of candidates has been chosen. In any case, it is paired with the selection procedure by the need of training the ML model on the augmented tables. To represent this coupling, we combine the two tasks in a single Merge step in Figure 1.